### 17 Mar 2026 v1

# Yet another characterization of Birkhoff Billiards inside discs

Abstract.
In this short paper, we show a characterization of Birkhoff Billiards inside discs which is related to the expansion of the formal Lazutkin conjugacy at the boundary.
2020 Mathematics Subject Classification:
1. Introduction and statement of our main results
Birkhoff billiards have been introduced almost a century ago by Birkhoff [3] as an example of a dynamical system that is simple to describe, but yet extremely rich in features. The system describes a particle moving inside an open domain $\Omega\subset\mathbb{R}^{2}$ along geodesics (straight lines) with the rule that, upon collisions with the boundary $\partial\Omega$, the particle reflects according to the law of optical reflection (i.e. the angle of incidence equals the angle of reflection). Despite the apparent simplicity of the problem, understanding billiard dynamics in an arbitrary domain is extremely challenging. Although much progress has recently taken place (see e.g. the survey [8]), many natural questions have remained open since their introduction and are very active topics of research.
In order to fix ideas, we assume $\partial\Omega$ to be smooth (i.e. parametrized by a $C^{\infty}$ curve) and strongly convex (i.e. such that the curvature at any point is strictly positive); we let $\ell$ denote the length of the boundary $\partial\Omega$ and $\mathbb{T}_{\ell}=\mathbb{R}/\ell\mathbb{Z}$. We denote with $(s,\varphi)\in\mathcal{M}=\mathbb{T}_{\ell}\times[0,\pi]$ the Birkhoff coordinates on the phase space, where $s$ is the arc-length parameter corresponding to a collision point and $\varphi$ is the angle that the outgoing trajectory forms with the positively oriented tangent to the boundary at that point. A particle following the trajectory identified by $(s,\varphi)$ will hit the boundary $\partial\Omega$ again at a unique point identified by the arc-length parameter $s^{\prime}$ and then reflects off the boundary with a certain angle of reflection $\varphi^{\prime}$. The map $f:(s,\varphi)\mapsto(s^{\prime},\varphi^{\prime})$ is called the Billiard Map; since the domain has a smooth boundary, then the billiard map is also smooth.
It is elementary to show that $f$ is diffeomorphism of $\mathbb{T}_{\ell}\times[0,\pi]$ onto itself: let $\mathcal{J}$ denote the diffeomorphism $(s,\varphi)\mapsto(s,\pi-\varphi)$. Observe that $\mathcal{J}$ is an involution (i.e. $\mathcal{J}\circ\mathcal{J}=\textrm{Id}$) and, moreover, that $f\circ\mathcal{J}\circ f\circ\mathcal{J}=\textrm{Id}$, which implies that $f^{-1}=\mathcal{J}\circ f\circ\mathcal{J}$ (the map $\mathcal{J}$ is called a reversor for $f$ as it reverses the direction of time). This observation shows that the dynamics of $f$ in a neighbourhood of the boundary curve $\{\varphi=0\}$ also determines the dynamics of $f$ in a neighbourhood of the other boundary curve $\{\varphi=\pi\}$; for this reason, when studying the behaviour of $f$ near the boundary, one can restrict to one of the two boundary curves. It is customary to choose the curve $\{\varphi=0\}$.
If $\Omega$ is a disc (of radius $r=\ell/2\pi$), then elementary geometrical considerations yield an explicit formula for $f$:
(1.1) $f:(s,\varphi)=(s+2r\varphi,\varphi).$
Lazutkin (see e.g. [10]) was the first to observe that if the domain is smooth, the billiard map can be conjugated in a neighbourhood of the boundary $\{\varphi=0\}$ to an arbitrarily good approximation of (1.1). There are many slightly different versions of this fundamental result; the version we propose below can be obtained by combining the arguments in [9, Appendix B] and [4, Appendix A]; see also [11, Lemma 14.6].
Lemma 0.1.
Assume $\partial\Omega$ is smooth, for any $k>0$ there exists a neighbourhood $U$ of $\{\varphi=0\}$ and a diffeomorphism $\Psi:U\to V$, where $V\subset\mathbb{T}\times[0,\varepsilon)$ is a neighbourhood of $\mathbb{T}\times\{0\}$ which conjugates $f$ with the normal form
(1.2) $(x,y)\mapsto(x+y,y)+O(y^{2k+2}).$
We call the above normal form Lazutkin normal form of order $k$. Moreover, we can write $\Psi=(X,Y)$, where
(1.3) $X(s,\varphi)$ $=\sum_{j=0}^{k-1}X_{j}(s)\varphi^{2j}$ $Y(s,\varphi)$ $=\sum_{j=0}^{k}Y_{j}(s)\varphi^{2j+1}.$
Finally, if we normalize $X$ so that $X(0,\varphi)=0$ for every $\varphi$, then the functions $X_{j}$ and $Y_{j}$ are uniquely determined for $j=0,\cdots,k-1$.
Remark 0.2.
Since a Lazutkin normal form of order $k$ is also a Lazutkin normal form of order $k^{\prime}$ for $k^{\prime}<k$, we conclude that the sequence $X_{j}$ and $Y_{j}$ can be uniquely determined from $\Omega$. Notice that $X_{0}$ is a toral diffeomorphism, whereas $X_{j}$ for $j>0$ and $Y_{j}$ for $j\geq 0$ are real-valued functions.
Observe that if $k=1$, the function $X(s,\varphi)=X_{0}(s)$ simply amounts to a reparametrization of the boundary, which is classically known as the Lazutkin parametrization and that we are now about to describe. Let $\rho_{\textrm{al}}(s)$ denote the radius of curvature of the domain at the point parametrized by arc-length $s$ and $\theta_{\textrm{al}}(s)$ denote the angle that the positively oriented tangent to $\partial\Omega$ forms with the horizontal positive semi-axis. Recall that, by definition, the radius of curvature at $s$ is the derivative of the tangent angle with respect to arc-length, that is:
(1.4) $\dot{\theta}_{al}(s)$ $=\rho_{\textrm{al}}^{-1}(s).$
Integrating the above along the domain yields the following elementary version of Gauss–Bonnet formula, which will be useful in the sequel:
(1.5) $\int_{0}^{\ell}\rho_{\textrm{al}}^{-1}(s)ds=2\pi.$
Lazutkin explicitly computed (in [10]) the reparametrization formula from arc-length as follows:
(1.6) $X_{0}(s)=C\int_{0}^{s}\rho_{\textrm{al}}^{-2/3}(s^{\prime})ds^{\prime},$
where $C$ is the normalizing factor:
(1.7) $C=\left[\int_{0}^{\ell}\rho_{\textrm{al}}^{-2/3}(s^{\prime})ds^{\prime}\right]^{-1}.$
From the Lazutkin formula (1.6) it is clear that $X_{0}$ is a linear rescaling if and only if $\Omega$ is a disc (since $X_{0}^{\prime}$ is a constant if and only if $\rho_{\textrm{al}}$ is a constant).
In this paper we perform a study of $X_{1}(s)$ that allows to prove the following characterization of Birkhoff billiards in a disc.
Main Theorem.
A smooth domain $\Omega$ is a disc if and only if $X_{1}\equiv 0$.
Remark 0.3.
In this paper (and hence the statement of our Main Theorem above ) we assume $C^{\infty}$ smoothness of the domain out of convenience, but the result indeed holds also in finite smoothness (e.g. for $C^{8}$ domains).
The formula (1.1) together with uniqueness of the conjugacy $\Psi$ immediately implies that if $\Omega$ is a disc, then $X_{1}\equiv 0$; this paper concerns the proof of the reverse implication.
1.1. Comments and ideas for future developments
Remark 0.4.
Our Main Theorem is one of the many dynamical characterizations of billiards in disc. For instance Bialy in [2] showed that the only domain whose phase space is completely foliated by non-contractible invariant curves (i.e. the billiard is globally integrable) is a disc. Discs are also characterized by the length of periodic orbits; for instance it is possible to show (see [14, 13, 7, 5]) that if the maximal length of a $n$-gon inscribed in $\Omega$ equals the length of the regular $n$-gon inscribed in a disc of circumference $\ell$, then $\Omega$ is necessarily a disc. In the recent preprint [1] it is shown that knowledge of the Mather’s $\beta$ function at only one rotation number may suffice to determine that the domain is a disc.
Remark 0.5.
Lemma 0.1 shows that it is always possible to formally (i.e. up to $O(\varphi^{\infty})$) conjugate the billiard map “at the boundary” of the domain to some universal normal form, which is given by
$(\xi,\eta)\mapsto(\xi+\eta,\eta).$
The term universal above refers to the fact that the normal form itself does not contain any information about the domain. It is rather the conjugacy map $\Psi$ that, through this procedure, absorbs all information about the domain. It makes therefore sense to ask how much information about the domain can be retrieved by a partial knowledge of the formal conjugacy, and this is the direction in which this paper is progressing.
Remark 0.6 (On Lazutkin conjugacy equivalence).
. On the one hand, in the remark above it is observed that, in the limit $k\to\infty$, the normal form is stripped of all information on $\Omega$. On the other hand, for any $k$, some information is still present in the structure of the remainder terms in (1.2). Lazutkin (see e.g. [11, (14.3)]) considers the following change of variables
(1.8) $X(s,\varphi)$ $=C\int_{0}^{s}\rho_{\textrm{al}}^{-2/3}(s^{\prime})ds^{\prime},$ $Y(s,\varphi)$ $=4C\rho_{\textrm{al}}^{1/3}(s)\sin({\varphi/2}).$
This variation on (1.3) has the desirable property of preserving a symplectic form. The billiard map in the above coordinates (1.8) admits the expansion (see e.g. [11, (14.4)])
(1.9) $\left\{\begin{array}[]{rcl}x^{+}&=&x+y+\tilde{\alpha}_{3}(x)y^{3}+\tilde{\alpha}_{4}(x)y^{4}+O(y^{5})\\ y^{+}&=&y+\tilde{\beta}_{4}(x)y^{4}+O(y^{5}).\end{array}\right$.
Fierobe recently announced in [6] the proof that two billiard tables with the same expansion (1.9) are necessarily homothetic (i.e. a rescaling of one another). It can be calculated that
$\tilde{\alpha}_{3}$ $=\frac{1}{27}\left(\frac{\rho^{\prime}}{\rho}\right)^{2}-\frac{1}{36}\frac{\rho^{\prime\prime}}{\rho}+\frac{1}{96C^{2}\rho^{2/3}},$
$\tilde{\alpha}_{4}$ $=-\frac{4}{135}\left(\frac{\rho^{\prime}}{\rho}\right)^{3}+\frac{11}{270}\frac{\rho^{\prime}\rho^{\prime\prime}}{\rho^{2}}-\frac{1}{90}\frac{\rho^{\prime\prime\prime}}{\rho}-\frac{\rho^{\prime}}{360C^{2}\rho^{5/3}},$
$\tilde{\beta}_{4}$ $=\frac{2}{135}\left(\frac{\rho^{\prime}}{\rho}\right)^{3}-\frac{11}{540}\frac{\rho^{\prime}\rho^{\prime\prime}}{\rho^{2}}+\frac{1}{180}\frac{\rho^{\prime\prime\prime}}{\rho}+\frac{\rho^{\prime}}{720C^{2}\rho^{5/3}}.$
Notice that $\tilde{\beta}_{4}=-\tilde{\alpha}_{4}/2$. From our discussion above, it is clear that there should be a relation between $\tilde{\alpha}_{3},\tilde{\alpha}_{4}$ and $\tilde{\beta}_{4}$ and the functions appearing in the expansion (1.3). For instance, modifying $Y_{3}(s)$ will arbitrarily affect $\tilde{\alpha}_{3}$ (for instance, Lemma 0.1 ensures that we could find $Y_{1}(s)$ which would set $\tilde{\alpha}_{3}=0$; see e.g. Remark 0.16 for further relations).
Fierobe subsequently realized that the proof in [6] is incorrect. The paper uses the formulae for the billiard map which can be found in [11, Page 145]; unfortunately such formulae are wrong.
The result presented in this paper would allow (see again Remark 0.16) to prove that the domain is a disc if and only if $\tilde{\alpha}_{4}=\tilde{\beta}_{4}=0$.
The result stated in [6] is still without a proof; we translate it below here in a language that is more in tune with the setting of the present paper.
Question: Does the function $X_{1}(s)$ determine the domain modulo homotheties? [^1] It is worth recalling the following much stronger conjecture, attributed to V. Guillemin: two billiard domains are homothetic if and only if their billiard maps are topologically conjugate.
Remark 0.7 (Distribution of collision points of periodic orbits).
In [4], the authors consider axially symmetric strictly convex domains with boundary of class $C^{r}$, $r\geq 8$. Without loss of generality, let us assume that the 0 of the parametrization is given by one of the intersections of $\partial\Omega$ with the symmetry axis. Rather elementary symmetry considerations allow to prove that there always exists a periodic orbit of any rotation number that has 0 as a collision point; such orbit is unique for sufficiently small rotation number. It is proven in [4] that if $(x^{k}_{q})$, $k=0,1\cdots,q-1$, are the coordinates of the periodic orbit of rotation number $1/q$ expressed in the Lazutkin parametrization (1.6), then
(1.10) $x^{k}_{q}=\frac{k}{q}+\frac{\alpha(k/q)}{q^{2}}+O(q^{-4})$
for some function $\alpha\in C^{r-4}(\mathbb{T})$. Inspecting the proof (presented in [4, Appendix A]) reveals that $X_{1}(s)=-\alpha(X_{0}(s))Y_{0}(s)^{2}$. Since $Y_{0}(s)=2C\rho^{1/3}(s)\neq 0$, the assumption in our Main Theorem is equivalent to say that $\alpha$ in (1.10) equals zero. In other terms: we know that for an arbitrary symmetric domain, collision points of periodic orbits are Lazutkin-equidistributed modulo a correction $\alpha/q^{2}+O(1/q^{4})$. Our Main Theorem asserts that the leading order ($1/q^{2}$) of this correction is 0 if and only if the domain is a disc. A version of this remark can be also formulated for domains that are not necessarily symmetric.
Remark 0.8.
Of course it is natural to ask similar questions for $X_{j}$ with $j>1$. We expect that it would still be possible to find differential relations involving $X_{1},X_{2},\cdots,X_{j}$ that would imply that $\Omega$ is a disc. It is unclear that such relations would be as simple as the one presented in this paper.
Acknowledgements
This paper originated as an undergraduate summer research project for AG and MP funded by a Mathematical and Computational Sciences Research Award at the University of Toronto Mississauga under the supervision of JDS. JDS acknowledges support of the NSERC Discovery grant, reference number RGPIN-2022-04188. JDS also wishes to thank Fabio Pusateri, for our always insightful exchanges of ideas, and Amir Vig, who suggested the reference [6]. JDS and KC are thankful to Corentin Fierobe for useful discussions about his paper. The algebraic computations presented in this paper have been carried out using the free open-source mathematics software system SageMath, which is released under the GNU General Public Licence GPLv2+.
2. The main calculation
The purpose of this section is to obtain a relation involving $X_{1}$ and $\rho_{\textrm{al}}$; this relation will be used in the next section to prove our Main Theorem. One possible way to approach this problem is by computing explicit formulas for the change of coordinates performed in [4, Lemma 5.1] or [9, Appendix B]. Here we propose an equivalent method, that is slightly more geometric in spirit and has the additional advantage that all the computations are performed in the same coordinates.
2.1. Angular functions
We will from now on assume that the arc-length parametrization of $\partial\Omega$ is smooth (see Remark 0.3). Also, we find convenient to identify $\mathbb{R}^{2}$ with $\mathbb{C}$: correspondingly, we will denote points on $\partial\Omega\subset\mathbb{C}$ with the symbol $z$. This being established, it is useful to observe that we can express the arc-length parametrization as:
(2.1) $z(s)=\int_{0}^{s}e^{i\theta_{\textrm{al}}(s^{\prime})}ds^{\prime},$
where we assume that the point identified by $s=0$ is at the origin and tangent to the positive horizontal direction (in particular $\Omega$ is contained in the upper half plane). We begin by providing some definitions:
Definition 0.9.
Let $(z_{-},z_{0},z_{+})$ identify a triple of points on $\partial\Omega$. Let $\varphi^{+}$ (resp. $\varphi^{-}$) denote the angle between $z_{0}z_{+}$ (resp. $z_{0}z_{-}$) with the positively (resp. negatively) oriented tangent direction at $z_{0}$ (see Figure 1). Let $\varepsilon>0$: we say that $(z_{-},z_{0},z_{+})$ identifies a billiard $\varepsilon$-pseudocollision if $|\varphi^{+}-\varphi^{-}|<\varepsilon$.
Notice that if $(z_{-},z_{0},z_{+})$ are three subsequent collision points of an actual billiard trajectory, then such triple is an $\varepsilon$-pseudocollision for every $\varepsilon>0$. Notice moreover that $(z_{-},z_{0},z_{+})$ is an $\varepsilon$-pseudocollision if and only if so is $(z_{+},z_{0},z_{-})$. Let us introduce the notation ${\mathbb{A}}_{\varepsilon}=\mathbb{T}\times(-\varepsilon,\varepsilon)$ and let ${\mathbb{A}}_{\varepsilon}^{*}={\mathbb{A}}_{\varepsilon}\setminus(\mathbb{T}\times\{0\})$.
Definition 0.10 (Angular function).
Let $k\geq 0$; a smooth function $Z\in C^{\infty}({\mathbb{A}}_{\varepsilon},\partial\Omega)$ is called an angular function of order $k$ if:
  1. (1)
    $Z(\cdot,\eta)=Z(\cdot,-\eta)$;
  2. (2)
    for any $\eta\in(-\varepsilon,\varepsilon)$, the map $Z(\cdot,\eta)$ is an orientation-preserving diffeomorphism and $Z(0,\eta)=Z(0,0)$;
  3. (3)
    for any $(\xi,\eta)\in{\mathbb{A}}_{\varepsilon}^{*}$, the triple
    $(Z(\xi-\eta,\eta),Z(\xi,\eta),Z(\xi+\eta,\eta))$
    identifies a billiard $O(\eta^{2k+2})$-pseudocollision (here the notation $O(\cdot)$ is intended as $\eta\to 0$).
$_{+}zz_{-}$$z_{0}$$\varphi^{+}$$\varphi^{-}$
Figure 1. The triple $(z_{-},z_{0},z_{+})$ identifies a billiard $\varepsilon$-pseudocollision if $|\varphi^{+}-\varphi^{-}|<\varepsilon$
Let $Z$ be an angular function and define $Z^{\pm}(\xi,\eta)=Z(\xi\pm\eta,\eta)$. For $(\xi,\eta)\in{\mathbb{A}}_{\varepsilon}^{*}$ we define the functions $\Phi^{\pm}(\xi,\eta)$ as follows: [^2] The choice to use arctangent rather than argument in the definition is deliberate, in order to obtain convenient symmetry properties. An alternative approach would be to use a projective version of the argument function (i.e. the argument modulo $\pi$ so that $\varphi$ and $\pi-\varphi$ would then be equivalent).
(2.2) $\Phi^{\pm}(\xi,\eta)$ $=\textrm{arctg}\frac{\textrm{Im}\,\delta z^{\pm}(\xi,\eta)}{\textrm{Re}\,\delta z^{\pm}(\xi,\eta)},\textrm{where }\delta z^{\pm}(\xi,\eta)=e^{-i\theta(Z(\xi,\eta))}\left[{Z^{\pm}(\xi,\eta)}-Z(\xi,\eta)\right],$
and $\theta(z)$ denotes the angle between the positively oriented tangent vector to $\partial\Omega$ at $z$ and the positive horizontal direction. Notice that, since we chose to preserve orientation in Definition 0.10, for positive $\eta$, the complex number $\delta z^{+}(\xi,\eta)$ (resp. $\delta z^{-}(\xi,\eta)$) belongs to the top-right (resp. top-left) quadrant. It is hence clear that, for positive $\eta$, the functions $\Phi^{\pm}({\xi,\eta})$ give the angles $\varphi^{\pm}$ corresponding to the triple $(Z(\xi-\eta,\eta),Z(\xi,\eta),Z(\xi+\eta,\eta))$. Observe that since $\partial\Omega$ is smooth, we can guarantee that $\theta(s)$ is also smooth. Notice that our definition implies $Z^{\pm}(\xi,-\eta)=Z^{\mp}({\xi,\eta})$, hence $\Phi^{\pm}(\xi,\eta)=-\Phi^{\mp}(\xi,-\eta)$. Moreover, $\lim_{\eta\to 0}\Phi^{\pm}(\xi,\eta)=0$; we thus extend $\Phi^{\pm}$ to the whole annulus ${\mathbb{A}}_{\varepsilon}$ by continuity, by setting $\Phi^{\pm}(\cdot,0)=0$.
We claim that, with this definition, $\Phi^{\pm}$ is a smooth function. In fact it is clear that $\Phi^{\pm}$ is smooth on ${\mathbb{A}}_{\varepsilon}^{*}$; it is then a simple exercise [^3] The exercise is similar to checking that if a smooth function $f(x)$ goes to 0 at $x=0$, then $f(x)/x$ (extended by continuity to be equal to $f^{\prime}(0)$ at 0) is also smooth. to show that $\Phi^{\pm}$ is also smooth at $\eta=0$.
The above discussion implies that $\Phi^{+}+\Phi^{-}$ is an odd function of $\eta$, whereas $\Phi^{+}-\Phi^{-}$ is an even function of $\eta$. This observation justifies the fact that we only consider even powers of $\eta$ in item (3) in Definition 0.10.
Remark 0.11.
The above discussion suggests a condition equivalent to item (3) in Definition 0.10. Let us consider the (partial) Taylor expansion at $\eta=0$:
(2.3) $\Phi^{\pm}(\xi,\eta)=\sum_{j=1}^{2k+1}\Phi^{\pm}_{j}(\xi){\eta^{j}}+O(\eta^{2k+2}).$
The symmetry relations discussed above imply that the following facts are equivalent
  • $\Phi^{+}_{2l}\equiv 0$ for $l\in\{1,\cdots,k\}$;
  • $\Phi^{-}_{2l}\equiv 0$ for $l\in\{1,\cdots,k\}$;
  • $Z$ is an angular function of order $k$.
The relations appearing in any of the first two items will provide (differential) equations that can be solved to obtain explicit expressions for the function $Z$.
2.2. Relation between angular functions and high-order Lazutkin coordinates
We now exhibit a relation between the function $Z$ and the Lazutkin change of variables $\Psi$: more in detail, we show that the existence of an angular function $Z$ of order $k$ is equivalent to existence of a change of coordinates $\Psi_{k}$ given in Lemma 0.1. We call such a change of coordinates Lazutkin change of coordinates of order $k$.
Lemma 0.12.
The following are equivalent:
  1. (1)
    there exists a Lazutkin change of coordinates of order $k$;
  2. (2)
    there exists a angular function $Z$ of order $k$ on some ${\mathbb{A}}_{\varepsilon}$.
[…]
Let us now prove that (2) $\Rightarrow$ (1). It is convenient to write $Z$ in coordinates: let $Z_{\textrm{al}}:{\mathbb{A}}_{\varepsilon}\to\mathbb{T}_{\ell}$ be the function mapping $(\xi,\eta)$ to the arc-length parametrization of $Z(\xi,\eta)$; without loss of generality we can assume that $Z_{\textrm{al}}(0,\eta)=0$. The argument used above item suggests that $Z_{\textrm{al}}$ should be related to the first component of the inverse of the conjugacy $\Psi$. Our strategy to prove indeed to construct a function $\Phi(\xi,\eta)$ in such a way that $(Z_{\textrm{al}}(\xi,\eta),\Phi(\xi,\eta))$ is a smooth diffeomorphism on ${\mathbb{A}}_{\varepsilon}$ and work from there. The definition of the function $\Phi$ it not unique; we propose a definition that manifestly preserves our symmetries. We take
$\Phi(\xi,\eta)=\frac{\Phi^{+}(\xi,\eta)+\Phi^{-}(\xi,\eta)}{2}.$
It was already observed that $\Phi$ is an odd function of $\eta$, i.e. $\Phi\circ J=-\Phi$; we now need to prove that $(\xi,\eta)\mapsto(Z_{\textrm{al}}(\xi,\eta),\Phi(\xi,\eta))$ is invertible in a suitable neighbourhood of $\{\eta=0\}$. Since $Z_{\textrm{al}}(\cdot,\eta)$ is an orientation-preserving diffeomorphism, we conclude that $\frac{\partial Z_{\textrm{al}}}{\partial\xi}>0$ everywhere. Moreover, since $Z_{\textrm{al}}(\cdot,\eta)=Z_{\textrm{al}}(\cdot,-\eta)$, we conclude that $\frac{\partial Z_{\textrm{al}}}{\partial\eta}|_{\eta=0}=0$. In order to conclude invertibility in some neighbourhood of $\{\eta=0\}$, it is then enough to show that $\frac{\partial\Phi}{\partial\eta}|_{\eta=0}\neq 0$; it is not difficult to check (see e.g. [11, (14.1b)]) that the following estimate holds everywhere (here $\rho_{\min}$ (resp. $\rho_{\max}$}) denotes the minimum (resp. maximum) of $\rho_{\textrm{al}}$):
$2\rho_{\min}|\Phi^{\pm}|\leq|Z_{\textrm{al}}^{\pm}-Z_{\textrm{al}}|\leq 2\rho_{\max}|\Phi^{\pm}|.$
Hence, if we can show that $\pm\frac{\partial Z_{\textrm{al}}^{\pm}}{\partial\eta}>0$, we can conclude that $\frac{\partial\Phi^{\pm}}{\partial\eta}>0$. By definition,
$\frac{\partial Z_{\textrm{al}}^{\pm}}{\partial\eta}=\pm\frac{\partial Z_{\textrm{al}}}{\partial\xi}+\frac{\partial Z_{\textrm{al}}}{\partial\eta}.$
If $\eta=0$, once again we have $\frac{\partial Z_{\textrm{al}}}{\partial\eta}=0$, which then proves our claim since $\frac{\partial Z_{\textrm{al}}}{\partial\xi}>0$ everywhere. The map $(Z_{\textrm{al}}(\xi,\eta),\Phi(\xi,\eta))$ is therefore a smooth orientation-preserving diffeomorphism in a neighbourhood of $\{\eta=0\}$. Its image contains a neighbourhood $\tilde{U}$ of $\{\varphi=0\}$ in phase space; let $\tilde{\Psi}$ be its inverse in such neighbourhood. Then $\tilde{\Psi}$ has the correct symmetries to be written in the form (1.3), but possibly it has extra terms in the expansion; we define $\Psi=(X,Y)$ to be the truncation of $\tilde{\Psi}$ of its Taylor expansion up to terms of order $\varphi^{2k}$ (for $X$) and order $\varphi^{2k+1}$ (for $Y$); the function $\Psi$ will be a diffeomorphism on a possibly smaller neighbourhood of $\{\varphi=0\}$ that we call $U$. Since $Z$ was an angular function, $\Phi(\xi,\eta)$ and $\Phi^{\pm}(\xi,\eta)$ differ by $O(\eta^{2k-2})$, which implies that $\Psi\circ f^{\pm 1}(s,\varphi)=(X(s,\varphi)\pm Y(s,\varphi),Y(s,\varphi))+O(\eta^{2k+2})$, proving that $\Psi$ is a high-order Lazutkin change of coordinates. ∎ If $Z$ is an angular function of order $k$, extracting the 2k12k-1-jet of $Z_{\textrm{al}}$ at $\eta=0$ we obtain, in particular, the following expansion:
$Z_{\textrm{al}}(\xi,\eta)$ $=\sum_{j=0}^{k-1}Z_{\textrm{al},j}(\xi)\eta^{2j}+O(\eta^{2k}).$
It is clear by Lemma 0.12 that there should be a relation between the functions $Z_{\textrm{al},j}$ and the functions $X_{j}$ and $Y_{j}$ given in (1.3). For instance, we immediately obtain:
(2.4) $Z_{\textrm{al},0}\circ X_{0}$ $=\text{Id}$ $Z_{\textrm{al},1}=-\frac{X_{1}}{\dot{X}_{0}Y_{0}^{2}}\circ Z_{\textrm{al},0}.$
In principle, one can obtain algebraic relations for arbitrary $k$ between the functions $Z_{\textrm{al},k}$ and $X_{j}$ and $Y_{j}$ for $0\leq j\leq k$. The immediate observation that stems from the above relation is that $X_{1}\equiv 0$ if and only if $Z_{\textrm{al},1}\equiv 0$. We can thus restate our Main Theorem as follows:
Theorem 0.13.
The domain $\Omega$ is a disc if and only if $Z_{\textrm{al},1}\equiv 0$.
In order to further simplify the computations that will follow, it turns out to be convenient to express $Z$ in Lazutkin coordinates. Let $Z:{\mathbb{A}}_{\varepsilon}\to\mathbb{T}$ be the function that maps $(\xi,\eta)$ to the Lazutkin coordinate corresponding to the point $Z(\xi,\eta)$. Of course we can write a new expansion:
$Z(\xi,\eta)$ $=\sum_{j=0}^{k-1}Z_{\El,j}(\xi)\eta^{2j}+O(\eta^{2k}).$
The advantage in the above expression is that (1.6) implies that $Z_{\El,0}=X_{0}\circ Z_{\textrm{al},0}=\text{Id}$, hence:
$Z(\xi,\eta)$ $=\xi+\sum_{j=1}^{k-1}Z_{\El,j}(\xi)\eta^{2j}+O(\eta^{2k}).$
Observe that $Z_{\textrm{al},1}\equiv 0$ if and only if $Z_{\El,1}\equiv 0$. According to Remark 0.11, in order to find an equation involving $Z_{\El,1}$, we need to set to zero the even terms of the expansion (2.3). This is the content of the following lemma. Let $\rho:\mathbb{T}\to\mathbb{R}_{>0}$ denote the radius of curvature of $\partial\Omega$ in the Lazutkin parametrization. We will denote with a prime differentiation with respect to the Lazutkin parameter (recall that a dot denoted differentiation with respect to arc-length). In order to simplify the notation we also denote by $A=Z_{\El,1}$. It is also useful to rewrite some formulas in the Lazutkin parametrization. It is a simple exercise to check that (2.1) can be written as (recall (1.6)):
(2.5) $z(x)=\frac{1}{C}\int_{0}^{x}e^{i\theta(x^{\prime})}\rho^{2/3}(x^{\prime})dx^{\prime},$
where $\theta$ is the angle $\theta$ expressed in Lazutkin parametrization; notice that the formula corresponding to (1.4) reads:
(2.6) $\theta^{\prime}=\frac{1}{C}\rho^{-1/3};$
in particular, the Gauss–Bonnet formula reads $\int_{\mathbb{T}}\rho^{-1/3}=2\pi C$.
Lemma 0.14.
Recall the expansion:
$\Phi^{+}(\xi,\eta)=\sum_{j=1}^{2k+1}\Phi_{j}^{+}(\xi)\eta^{j}+o(\eta^{k}).$
Then
$\Phi^{+}_{2}$ $=0$
$\Phi^{+}_{4}$ $=\frac{1}{4C\rho^{1/3}}\cdot\left(A^{\prime\prime}+\frac{8}{135}\left(\frac{\rho^{\prime}}{\rho}\right)^{3}-\frac{11}{135}\frac{\rho^{\prime}\rho^{\prime\prime}}{\rho^{2}}+\frac{1}{45}\frac{\rho^{\prime\prime\prime}}{\rho}+\frac{1}{180C^{2}}\frac{\rho^{\prime}}{\rho^{5/3}}\right),$
where all the functions are evaluated in (and differentiated with respect to) the variable $\xi$.
Let us recall the definition of $\Phi^{+}$ from (2.2); we now rewrite the complex vector $\delta z^{\pm}$ using (2.5) and (2.6) obtaining:
(2.7) $\delta z^{\pm}(\xi,\eta)=\frac{e^{-i\theta(Z(\xi,\eta))}}{C}\int_{Z(\xi,\eta)}^{Z^{\pm}(\xi,\eta)}e^{i\theta(x)}\rho^{2/3}(x)dx.$
Since $Z_{\El,0}=\text{Id}$ and we are in Lazutkin parametrization, we conclude that $Z_{\El,0}$ is an angular function of order $1$, hence $\Phi^{+}_{2}=0$ (this follows also from an explicit computation).
Since moreover arctg is an odd function of its argument, it holds that $\Phi_{4}^{+}$ equals the fourth Taylor coefficient of $\textrm{tg}\Phi^{+}$; in other words, we can avoid taking the arctg in order to compute the required coefficient.
Applying Lemma 0.21 we can thus write:
$\textrm{tg}\Phi^{+}(\xi,\eta)=\frac{S_{2}\eta^{2}+S_{3}\eta^{3}+S_{4}\eta^{4}+S_{5}\eta^{5}}{C_{1}\eta+C_{2}\eta^{2}+C_{3}\eta^{3}+C_{4}\eta^{4}}+O(\eta^{5})$
The fourth-order term in this expansion is thus given by (see once again Lemma 0.21 for the definitions of the functions $\tilde{S}_{j}$ and $\tilde{C}_{j}$):
$\Phi^{+}_{4}$ $=C^{-1}\rho^{-1/3}\cdot\left(\tilde{S}_{5}-\tilde{C}_{2}\tilde{S}_{4}+(\tilde{C}_{2}^{2}-\tilde{C}_{3})\tilde{S}_{3}-(\tilde{C}_{2}^{3}-2\tilde{C}_{2}\tilde{C}_{3}+\tilde{C}_{4})\tilde{S}_{1}\right).$
Carrying out the algebra yields the desired formula. ∎ We conclude that, in order for $Z$ to be an angular function of order $2$, the function $A=Z_{\El,1}$ solves the following differential equation:
(2.8) $A^{\prime\prime}+\frac{8}{135}\left(\frac{\rho^{\prime}}{\rho^{3}}\right)^{3}-\frac{11}{135}\frac{\rho^{\prime}\rho^{\prime\prime}}{\rho^{2}}+\frac{1}{45}\frac{\rho^{\prime\prime\prime}}{\rho}+\frac{1}{180C^{2}}\frac{\rho^{\prime}}{\rho^{5/3}}=0.$
with periodic boundary conditions and (by our choice of normalization) $A(0)=0$. We will study solutions the differential equation in the next section.
Remark 0.15.
Among the four terms involving $\rho$, three have differential degree $3$ and the remaining term has differential degree $1$ (once renormalized with $C^{2}\rho^{2/3}$). Terms with different differential degrees come from different terms in the expansion of the complex exponential function in (2.7).
Remark 0.16.
Observe that (2.8) is equivalent to the equation
$\frac{1}{4}A^{\prime\prime}+\tilde{\beta}_{4}=0,$
where $\tilde{\beta}_{4}$ is defined in Remark 0.6. This is of course to be expected, since $\tilde{\beta}_{4}$ is the correction of order $4$ to the angular component $y$ of the billiard map in classical Lazutkin coordinates, and $A$ is exactly designed to compensate such correction.
3. Proof of the Main Theorem
In this section we prove the Main Theorem; we begin by some elementary manipulations
3.1. Solving the differential equation (2.8)
Performing the convenient substitution $h=\rho^{-1/3}/2\sqrt{2}$ in (2.8), we obtain:
$A^{\prime\prime}$ $=\frac{1}{15}\frac{h^{\prime\prime\prime}}{h}-\frac{1}{15}\frac{h^{\prime}h^{\prime\prime}}{h^{2}}+\frac{2}{15C^{2}}hh^{\prime}=$
(3.1) $=\frac{1}{15}\left[\frac{h^{\prime\prime}}{h}+\frac{h^{2}}{C^{2}}\right]^{\prime}.$
Observe that, with this substitution, the Gauss–Bonnet formula (1.5) amounts to a normalization condition for $h$ that we write as follows:
(3.2) $2\sqrt{2}\int_{\mathbb{T}}hd\xi$ $=2\pi C.$
Integrating once the above differential equation, we therefore obtain
(3.3) $A^{\prime}=\frac{1}{15}\left[\frac{h^{\prime\prime}}{h}+\frac{h^{2}}{C^{2}}-C_{\textrm{BV}}\right],$
where $C_{\textrm{BV}}$ is an integration constant that can be fixed in such a way to ensure periodicity of $A$ — that is, we set $\int A^{\prime}=0$, which yields, after integrating by parts, the manifestly positive quantity:
(3.4) $C_{\textrm{BV}}$ $=\int_{\mathbb{T}}\left[\left(\frac{h^{\prime}}{h}\right)^{2}+\frac{h^{2}}{C^{2}}\right]d\xi.$
Thus, integrating one more time and imposing the condition $A(0)=0$, we conclude:
(3.5) $15\,A(\xi)=\int_{0}^{\xi}\frac{h^{\prime\prime}(\xi^{\prime})}{h(\xi^{\prime})}d\xi^{\prime}+\frac{1}{C^{2}}\int_{0}^{\xi}h(\xi^{\prime})^{2}d\xi^{\prime}-C_{\textrm{BV}}\xi,$
Remark 0.17.
It may be useful for future work to explicitly write the expression (3.3) in terms of the curvature function: here derivatives are with respect to Lazutkin coordinates:
$A^{\prime}=\frac{4}{135}\left(\frac{\rho^{\prime}}{\rho}\right)^{2}-\frac{1}{45}\frac{\rho^{\prime\prime}}{\rho}+\frac{1}{120C\rho^{2/3}}-\frac{C_{\textrm{BV}}}{15}.$
Remark 0.18.
It is instructive to observe that the quantity $C_{\textrm{BV}}$ has some intrinsic meaning. Reverting in (3.4) the substitution for $\rho$ yields
$C_{\textrm{BV}}=\frac{1}{72}\int_{\mathbb{T}}8\left(\frac{\rho^{\prime}}{\rho}\right)^{2}+\frac{9}{C^{2}}\rho^{-2/3}d\xi;$
changing variables to arc-length and recalling (1.7) gives:
$C_{\textrm{BV}}=\frac{1}{72}\int_{\mathbb{T}_{\ell}}\rho_{\textrm{al}}^{-2/3}ds\cdot\int_{\mathbb{T}_{\ell}}\frac{8\dot{\rho}_{\textrm{al}}^{2}+9}{\rho_{\textrm{al}}^{4/3}}ds.$
The two integrals appearing in the above expression are (multiples of) two of the so-called Marvizi–Melrose invariants $\mathcal{I}_{1}$ (which we called $C^{-1}$) and $\mathcal{I}_{2}$ (see e.g. [12, (4.6)], where they are defined in terms of the curvature $\kappa=\rho_{\textrm{al}}^{-1}$ or [15, Section 2.5]). Such quantities are Laplace-spectral and Marked-Length-Spectral invariants.
This perhaps surprising relation is of course not accidental. As remarked in Remark 0.7, the function $A$ appears (in some form) in the asymptotic distribution of collision points of periodic orbits; this in turn is tightly related with the asymptotics of the length of periodic orbits of small rotation number –or, equivalently, the coefficients of the Mather’s $\beta$ function at 0– and it is well known (see once again [12, Theorem 5.15] and [15]) that such coefficients are algebraically related with the Marvizi–Melrose invariants.
3.2. Characterizing the disc
We now assume that $A\equiv 0$ and write an equation for $h$; substituting $A^{\prime\prime}\equiv 0$ in (3.1) we obtain the equation:
$\left[\frac{h^{\prime\prime}}{h}+\frac{h^{2}}{C^{2}}\right]^{\prime}=0.$
Observe that the constant $C$ appearing in the above equation itself depends on the solution $h$ via the relation (3.2); since $C$ amounts to a normalization, this complication can be avoided by writing the equation in terms of the unrenormalized Lazutkin parameter. In other terms, performing the substitution $t=C\xi$ (and denoting the derivative with respect to $t$ with a subscript) we obtain:
$C^{-3}\left[\frac{h_{tt}}{h}+h^{2}\right]_{t}=0,$
in which the dependence on $C$ can be factored out. We now integrate once with respect to $t$ and multiply by $h$ (since $h\neq 0$) obtaining:
(3.6) $h_{tt}=-h^{3}+C_{\star}h,$
where $C_{\star}$ is a constant of integration. Any positive solution of (3.6) (depending on the parameter $C_{\star}$) identifies a curvature function $\rho(t)=h(t)^{-3}/8$; in turn any curvature function identifies a planar curve $\Gamma(t)$ by (the unrenormalized versions of) (2.5) and (2.6):
$\Gamma(t)=\int_{0}^{t}e^{i\int_{0}^{t^{\prime}}\rho^{-1/3}(t^{\prime\prime})dt^{\prime\prime}}\rho^{2/3}(t^{\prime})dt^{\prime}.$
We will declare a solution $h$ to be admissible if $\Gamma$ is a Jordan curve (i.e. the boundary of a domain). In particular both $\Gamma$ and $h$ need to be periodic with the same period (although the primitive period of $\Gamma$ may be a multiple of the primitive period of $h$). Additionally $h$ needs to satisfy the Gauss–Bonnet formula:
(3.7) $2\sqrt{2}\oint hdt=2\pi,$
where $\oint$ denotes the integral on the period of $\Gamma$. [^4] Of course the mere periodicity of $h$ and (3.7) are only necessary for $\Gamma$ to be Jordan but not sufficient. We will conclude this section with the proof of the following lemma, from which our Main Theorem ensues.
Lemma 0.19.
For any $C_{\star}>0$, the differential equation (3.6) has a unique admissible periodic solution: this solution is given by the constant $h\equiv\sqrt{C_{\star}}$.
First of all, it is obvious to check that the constant solution $h\equiv\sqrt{C_{\star}}$ satisfies (3.6). In that case $\Gamma$ is a circle (of radius $C_{\star}^{-3/2}/8$) and $h$ is certainly admissible.
Before continuing with the proof, it is first convenient to normalize (3.6) one last time, so that $C_{\star}$ does not appear in the equation. Let $h=\sqrt{C_{\star}}\zeta$ and $\tau=\sqrt{C_{\star}}t$ so that (3.6) reads:
(3.8) $\zeta_{\tau\tau}=-\zeta^{3}+\zeta.$
Observe that with this substitution, the Gauss–Bonnet constraint (3.7) still reads
(3.9) $2\sqrt{2}\oint\zeta d\tau=2\pi.$
The proof of the lemma will be complete if we can show that any non-constant periodic solution $\zeta$ of (3.8) fails to satisfy (3.9). Let us consider the following function of $\zeta$ and $\zeta_{\tau}$:
$\mathcal{H}(\zeta,\zeta_{\tau}):=\frac{1}{2}\zeta_{\tau}^{2}+\frac{\zeta^{4}}{4}-\frac{\zeta^{2}}{2}+\frac{1}{4}.$
An elementary computation shows that $\mathcal{H}$ is a conserved quantity (in fact, $\mathcal{H}$ is a Hamiltonian function for the system given by (3.6)). We thus study the level sets $\mathcal{H}=E$, which correspond to possible trajectories of (3.6). It is elementary to check that such level sets are symmetric with respect to the transformation $\zeta\mapsto-\zeta$; moreover the system admits trajectories for any $E\geq 0$; such trajectories are always bounded. If, moreover, $E<1/4$, such trajectories do not traverse $\zeta=0$. Observe moreover that $E=0$ corresponds to the constant solution $\zeta\equiv 1$ discussed earlier; since we can only admit positive solutions, we will therefore assume $E\in(0,1/4)$.
We can thus solve for $\zeta_{\tau}$, obtaining:
$\zeta_{\tau}=\pm\sqrt{2E-\frac{1}{2}-\frac{\zeta^{4}}{2}+\zeta^{2}}=\pm\sqrt{2E-\frac{(\zeta^{2}-1)^{2}}{2}}.$
This shows that the integral curves of (3.6) in the plane $(\zeta,\zeta_{\tau})$ are closed and symmetric with respect to the axis $\zeta_{\tau}=0$. The trajectory with energy $E$ oscillates between the two positive solutions of the bi-quadratic equation $4E=(\zeta^{2}-1)^{2}$, which satisfy the relation
$\zeta^{2}_{\pm}=1\pm 2\sqrt{E}.$
Any fixed $E\in(0,1/4)$ corresponds to a trajectory that is periodically oscillating back and forth between $\zeta_{-}(E)$ and $\zeta_{+}(E)$; let us denote its prime period by $T(E)$.
For given $E$, let us now define the integral
$\Theta(E)=2\sqrt{2}\int_{0}^{T(E)}\zeta d\tau.$
As observed earlier, this quantity measures the angle between the vectors $\Gamma_{\tau}(0)$ and $\Gamma_{\tau}(T(E))$, tangent to the curve $\Gamma$ respectively at $\tau=0$ and $\tau=T(E)$. The function $\Theta(E)$ can be computed explicitly (the crucial factor 2 accounts for the fact that, in one period, the function oscillates from $\zeta_{-}$ to $\zeta_{+}$ and back):
$\Theta(E)=2\cdot 2\sqrt{2}\int_{\zeta_{-}}^{\zeta_{+}}\frac{\zeta}{\sqrt{2E-\frac{1}{2}(\zeta^{2}-1)^{2}}}d\zeta.$
Performing the substitutions $u=\zeta^{2}-1$ and $v=u/(2\sqrt{E})$ yields:
$\Theta(E)=2\sqrt{2}\int_{-2\sqrt{E}}^{2\sqrt{E}}\frac{du}{\sqrt{2E-\frac{1}{2}u^{2}}}=4\int_{-1}^{1}\frac{dv}{\sqrt{1-v^{2}}}=4\pi.$
Remarkably, the quantity $\Theta(E)$ is independent of $E$, and moreover it is twice $2\pi$. As observed earlier, the period of $\Gamma$ is necessarily a multiple of the period $T(E)$ of $\zeta$: we conclude that the left hand side of (3.9) equals a positive multiple of $4\pi$, hence it never satisfies the equation. We conclude that no non-constant periodic solution is admissible. ∎
Remark 0.20.
In the above proof we show that non-constant periodic solutions of (2.8) with $A^{\prime\prime}=0$ are curvature functions of curves that necessarily self-intersect. In fact, there are additional considerations that would rule the solutions inadmissible. For instance it would possible be to show that for any $E\in(0,1/4)$ the curves $\Gamma$ corresponding to the value $E$ are not closed (see Figure 2 below for some numerical solutions)
Figure 2. Curves $\Gamma$ corresponding to a full period of the solution $\zeta$ with $E=0.005$, $E=0.02$ and $E=0.05$ respectively (left to right). The curves appear not to be periodic, and to complete an angle of $4\pi$.
4. Computing $X_{1}$
Since $A=Z_{\El,1}$, formulae (2.4) and (3.5) can be exploited to find $X_{1}$. Indeed, let us recall that
$Z_{\El,1}(\xi)=A(\xi)=\frac{1}{15}\int_{0}^{\xi}\frac{h^{\prime\prime}(\xi^{\prime})}{h(\xi^{\prime})}d\xi^{\prime}+\frac{1}{15C^{2}}\int_{0}^{\xi}h(\xi^{\prime})^{2}d\xi^{\prime}-\frac{1}{15}C_{\textrm{BV}}\xi,$
where all derivatives are with respect to Lazutkin. First let us substitute $\xi^{\prime}=X_{0}(s^{\prime})=C\int_{0}^{s^{\prime}}\rho_{\textrm{al}}^{-2/3}(\zeta)d\zeta$, $d\xi^{\prime}=C\rho_{\textrm{al}}^{-2/3}(s^{\prime})ds^{\prime}$ to obtain
$Z_{\textrm{al},1}(X_{0}(s))=\frac{1}{15}\int_{0}^{s}C\rho_{\textrm{al}}^{-2/3}(s^{\prime})\frac{h^{\prime\prime}(X_{0}(s^{\prime}))}{h(X_{0}(s^{\prime}))}ds^{\prime}$
$+\frac{1}{15C}\int_{0}^{s}\rho_{\textrm{al}}^{-2/3}(s^{\prime})h(X_{0}(s^{\prime}))^{2}ds^{\prime}-\frac{1}{15}C_{\textrm{BV}}X_{0}(s),$
where $s$ is such that $X_{0}(s)=\xi$, and put $h=\frac{\rho^{-1/3}}{2\sqrt{2}}$ to obtain
$Z_{\textrm{al},1}(X_{0}(s))=\frac{4C}{135}\int_{0}^{s}\frac{\rho^{\prime}(X_{0}(s^{\prime}))^{2}}{\rho_{\textrm{al}}(s^{\prime})^{8/3}}ds^{\prime}-\frac{C}{45}\int_{0}^{s}\frac{\rho^{\prime\prime}(X_{0}(s^{\prime}))}{\rho_{\textrm{al}}(s^{\prime})^{5/3}}ds^{\prime}$
$+\frac{1}{120C}\int_{0}^{s}\frac{ds^{\prime}}{\rho_{\textrm{al}}(s^{\prime})^{4/3}}-\frac{C_{\textrm{BV}}}{15}X_{0}(s).$
It can be easily find that
$\rho^{\prime}(X_{0}(s))=\frac{\rho_{\textrm{al}}^{2/3}(s)}{C}\dot{\rho}_{\textrm{al}}(s),$
$\rho^{\prime\prime}(X_{0}(s))=\frac{1}{C^{2}}\rho_{\textrm{al}}^{4/3}(s)\ddot{\rho}_{\textrm{al}}(s)+\frac{2}{3C^{2}}\rho_{\textrm{al}}^{1/3}(s)\dot{\rho}_{\textrm{al}}(s)^{2},$
where dots represent differentiation with respect to arc-length. Thus
$Z_{\textrm{al},1}(X_{0}(s))=\frac{2}{135}\int_{0}^{s}\frac{\dot{\rho}_{\textrm{al}}(s^{\prime})^{2}}{C\rho_{\textrm{al}}(s^{\prime})^{4/3}}ds^{\prime}-\frac{1}{45}\int_{0}^{s}\frac{\ddot{\rho}_{\textrm{al}}(s^{\prime})}{C\rho_{\textrm{al}}(s^{\prime})^{1/3}}ds^{\prime}$
$+\frac{1}{120C}\int_{0}^{s}\frac{ds^{\prime}}{\rho_{\textrm{al}}(s^{\prime})^{4/3}}-\frac{C_{\textrm{BV}}}{15}X_{0}(s).$
By (2.4) we have $X_{1}(s)=-Z_{\textrm{al},1}(X_{0}(s))\dot{X}_{0}(s)Y_{0}^{2}(s)$, thus finally we obtain
$X_{1}(s)=-\int_{0}^{s}\bigg[\frac{8C^{2}\dot{\rho}_{\textrm{al}}(s^{\prime})^{2}}{135\rho_{\textrm{al}}(s^{\prime})^{4/3}}-\frac{{4C^{2}\ddot{\rho}_{\textrm{al}}(s^{\prime})}}{45\rho_{\textrm{al}}(s^{\prime})^{1/3}}+\frac{C^{2}}{30\rho_{\textrm{al}}(s^{\prime})^{4/3}}-\frac{4C^{3}C_{\textrm{BV}}}{15\rho_{\textrm{al}}^{2/3}(s^{\prime})}\bigg]ds^{\prime}.$
Appendix 0.A Derivatives
Lemma 0.21.
Let $\Omega$ be an arbitrary convex domain with a $C^{8}$ boundary, $x$ denote the Lazutkin parametrization of its boundary. For $x\in\mathbb{T}$, recall that $\rho(x)$ denotes the radius of curvature of the boundary at the point on $\partial\Omega$ parametrized by $x$ and $\theta(x)$ the angle between the tangent at the same point and the positive horizontal semi-axis. Further, let $A\in C^{3}(\mathbb{T})$ and
$x_{0}=\xi+A(\xi)\eta^{2},\quad x_{+}=\xi+\eta+{A(\xi+\eta)}\eta^{2}.$
Then there exist functions $S_{j}$ and $C_{j}$ for $j=0,\cdots,5$ so that:
(0.A.1) $\int_{x_{0}}^{x_{+}}e^{i(\theta(x^{\prime})-\theta(x_{0}))}\rho(x^{\prime})^{2/3}dx^{\prime}=\sum_{j=0}^{5}C_{j}\eta^{j}+i\sum_{j=0}^{5}S_{j}\eta^{j}+O(\eta^{6})$
and the following formulas hold: let us renormalize
$S_{j}$ $=\frac{\rho^{1/3}}{C^{2}}\tilde{S}_{j}$ $C_{j}$ $=\frac{\rho^{2/3}}{C}\tilde{C}_{j}$
Then we have:
$\tilde{S}_{0}$ $=\tilde{S}_{1}=0$ $\tilde{S}_{2}$ $=\frac{1}{2}$ $\tilde{S}_{3}$ $=\frac{1}{6}\frac{\rho^{\prime}}{\rho}$
$\tilde{S}_{4}$ $=A^{\prime}-\frac{1}{27}\left(\frac{\rho^{\prime}}{\rho}\right)^{2}+\frac{5}{27}\frac{\rho^{\prime\prime}}{\rho}-\frac{1}{24C^{2}}\rho^{-2/3}.$
$\tilde{S}_{5}$ $=\frac{1}{2}A^{\prime\prime}+\frac{1}{3}A^{\prime}\frac{\rho^{\prime}}{\rho}+\frac{2}{135}\left(\frac{\rho^{\prime}}{\rho}\right)^{3}-\frac{4}{135}\frac{\rho^{\prime}\rho^{\prime\prime}}{\rho^{2}}+\frac{7}{360}\frac{\rho^{\prime\prime\prime}}{\rho}-\frac{1}{180C^{2}}\frac{\rho^{\prime}}{\rho^{5/3}},$
and
$\tilde{C}_{0}$ $=0$ $\tilde{C}_{1}$ $=1$ $\tilde{C}_{2}=\frac{1}{3}\frac{\rho^{\prime}}{\rho}$
$\tilde{C}_{3}=A^{\prime}-\frac{1}{27}\left(\frac{\rho^{\prime}}{\rho}\right)^{2}+\frac{1}{9}\frac{\rho^{\prime\prime}}{\rho}-\frac{1}{6C^{2}}\rho^{-2/3}$
$\tilde{C}_{4}$ $=\frac{1}{2}A^{\prime\prime}+\frac{1}{3}\frac{\rho^{\prime}}{\rho}A^{\prime}+\frac{1}{81}\left(\frac{\rho^{\prime}}{\rho}\right)^{3}-\frac{1}{36}\frac{\rho^{\prime}\rho^{\prime\prime}}{\rho^{2}}+\frac{1}{36}\frac{\rho^{\prime\prime\prime}}{\rho}-\frac{1}{24C^{2}}\frac{\rho^{\prime}}{\rho^{5/3}}.$
Integrating (2.6) yields:
$\theta(x^{\prime})-\theta(x_{0})=\int_{x_{0}}^{x^{\prime}}\theta^{\prime}(x^{\prime\prime})dx^{\prime\prime}=C^{-1}\int_{x_{0}}^{x^{\prime}}\rho^{-1/3}(x^{\prime\prime})dx^{\prime\prime}.$
We can thus write (0.A.1) as:
$\int_{x_{0}}^{x_{+}}e^{i(\theta(x^{\prime})-\theta(x_{0}))}\rho(x^{\prime})^{2/3}dx^{\prime}$ $=\int_{x_{0}}^{x_{+}}\exp\left(\frac{i}{C}\int_{x_{0}}^{x^{\prime}}\rho^{-1/3}(x^{\prime\prime})dx^{\prime\prime}\right)\rho(x^{\prime})^{2/3}dx^{\prime}.$
The result then follows by performing a formal Taylor expansion in $\eta$ around $\eta=0$ and carrying out the algebra. ∎